If the roots of the equation  (a-b) x2 + (b-c) x + (c-a) = 0  are 
equal , then prove that  2a = b + c .

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
As we see that x = 1 satisfy this
Hence both roots are 1.
 
Hence
Product of roots = 1 = c-a /a-b
 
a-b = c -a 
 
2a = b +c